OpenStudy (anonymous):

Simplify e^6f^3/e^4f^9 x f^6 e^2/e^6 e^2f^6/f^-6 e^2

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@ParthKohli

OpenStudy (muzzack):

Final result : \[\LARGE\ e^2f ^{18} x\]

OpenStudy (anonymous):

o_o?

OpenStudy (igreen):

Nevermind.

OpenStudy (igreen):

Is this it? \(\Large\frac{e^6f^3}{e^4f^9} \times f^6\)

OpenStudy (anonymous):

Yep

OpenStudy (igreen):

\(e^2\) is your answer.

OpenStudy (anonymous):

simplify 3^-3

OpenStudy (igreen):

All the f's become \(f^0\) which is 0.

OpenStudy (igreen):

0.037

OpenStudy (anonymous):

1/9, -27, or 1/27

OpenStudy (igreen):

1/27 = 0.037037037 Exact same thing.

OpenStudy (anonymous):

Made a 100 on that one Preciate it! Talk to you tomorrow!

OpenStudy (igreen):

Bye!